8. Use the steam tables to calculate the enthalpy change (in kJ) of 2 kg-mol of water/steam when heated from:

a) 293 K & 100 kPa to 303 K & 100 kPa.

b) 400 K & 100 kPa to 900 K & 100 kPa..

 


1
Expert's answer
2022-01-07T09:34:02-0500

a) ∆H = m x Cp x (T2 - T1) = 2 kg-mol x 75.38 J/mol K x (303 - 293 K) = 1.5 kJ

b) ∆H = m x Cp x (T2 - T1) = 2 kg-mol x 36.57 J/mol K x (900 - 400 K) = 36.6 kJ